OpenStudy (anonymous):

Find the equation of parabola with Focus (-1,3) and P=4, axis parallel to x-asis.

OpenStudy (anonymous):

need help please!!!

OpenStudy (anonymous):

a) (y + 3)^ 2 = 16(x - 5) b) (y - 3)^2 = 4(x + 5) c) (y - 1)^2 = 4x d) (y +1 )^2 = 16(x + 1)

OpenStudy (anonymous):

help help help help

OpenStudy (anonymous):

a

OpenStudy (anonymous):

are u sure ??

OpenStudy (anonymous):

thanks brother

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

but i think Focus is (1,-3).....check ur question again

OpenStudy (anonymous):

the focus is (-1,3).. Can you show me the steps???

OpenStudy (anonymous):

or ur ans should be a) (y - 3)^ 2 = 16(x + 5)

OpenStudy (anonymous):

this parabola standard equation is (y-k)^2=4P(x-h) Focus (h+p,k)

OpenStudy (anonymous):

Ok,, I got it,, thanks.. I have another Questions..

OpenStudy (anonymous):

Determine the equation for a cycle tangent to both axes and the lines X=4 and Y=4

OpenStudy (anonymous):

(x-2)^2+(y-2)^2=4